OpenStudy (anonymous):

n a recent golf match, Tiger’s score was 4 less than Phil’s score. Their combined scores totaled 140. Which of the following pairs of equations can be used to determine their scores? Let p represent Phil’s score and t represent Tiger’s score. p + 140 = t p = 4 p + t = 140 p + t = 4 p + t = 140 p – t = 4 p – t = 140 p + t = 4

OpenStudy (kc_kennylau):

Let's break the question into two parts: (i) Tiger’s score was 4 less than Phil’s score. (ii) Their combined scores totaled 140.

OpenStudy (kc_kennylau):

Which part do you not understand? :)

OpenStudy (anonymous):

None of it, math is confusing

OpenStudy (kc_kennylau):

If you say "a" is 4 less than "b", what am I saying?

OpenStudy (kc_kennylau):

If I say*

OpenStudy (anonymous):

a-4=b?

OpenStudy (kc_kennylau):

exactly :)

OpenStudy (kc_kennylau):

now rearrange the equation so that 4 is on the right hand side :)

OpenStudy (anonymous):

a-b=4?

OpenStudy (kc_kennylau):

exactly, now apply that to (i)

OpenStudy (anonymous):

The answer is D right?

OpenStudy (kc_kennylau):

nope

OpenStudy (anonymous):

B?

OpenStudy (kc_kennylau):

nope, and stop guessing.

OpenStudy (anonymous):

well i tried to plug in the numbers like in the practice equation you gave me and those were the outcomes :/ am i doing something wrong?

OpenStudy (kc_kennylau):

Well, the equation I gave you was a-b=4

OpenStudy (kc_kennylau):

But both of your choices were p+t=4

OpenStudy (kc_kennylau):

which is + not -

OpenStudy (anonymous):

oh well let me try it again :)

OpenStudy (kc_kennylau):

ok :)

OpenStudy (anonymous):

p + t = 140 p – t = 4

OpenStudy (kc_kennylau):

bingo :D

OpenStudy (anonymous):

:D